# SEGAN: semi-supervised learning approach for missing data imputation

### v2

@@ Abstract
In many practical real-world applications, data missing is a very common phenomenon, making the development of data-driven artificial intelligence theory and technology increasingly difficult. Data completion is an important method for missing data preprocessing. Most existing missing data completion models directly use the known information in the missing data set but ignore the impact of the data label information contained in the data set on the missing data completion model. To this end, this paper proposes a missing data completion model SEGAN based on semi-supervised learning, which mainly includes three important modules: generator, discriminator and classifier. In the SEGAN model, the classifier enables the generator to make more full use of known data and its label information when predicting missing data values. In addition, the SEGAN model introduces a missing hint matrix to allow the discriminator to more effectively distinguish between known data and data filled by the generator. This paper theoretically proves that the SEGAN model that introduces a classifier and a missing hint matrix can learn the real known data distribution characteristics when reaching Nash equilibrium. Finally, a large number of experiments were conducted in this article, and the experimental results show that compared with the current state-of-the-art multivariate data completion method, the performance of the SEGAN model is improved by more than 3%.
@@ Introduction
In the information age of today, the issue of data quality has become increasingly pivotal, with the problem of missing data being particularly widespread. This issue is not confined to any single sector but is prevalent across various industries such as healthcare, meteorology, and advertising, and it has a significant impact on data analysis and decision-making. Missing data can lead to incompleteness of information and may also cause biases in data analysis, thereby affecting the accuracy and reliability of decision-making. Faced with this challenge, the academic and industrial communities have devoted efforts researching data imputation techniques, aiming to find more effective methods to address this issue [bib1]. Historically, researchers have developed a variety of data imputation techniques to address the issue of missing data. Early methods often relied on statistics, for instance filling in missing values by inserting mean, median, or mode [bib2] [bib3]. While these methods are straightforward in computation, they overlook the potential correlations and dynamic changes among data, often failing to achieve the desired imputation result. With the development of machine learning and deep learning methods, parametric imputation algorithms based on these methods have emerged [bib1]. These algorithms predict missing values by learning the complex relationships between data, demonstrating more accurate and efficient imputation capabilities. Moreover, the information contained in data labels is crucial for research on data imputation and data analysis tasks [bib37]  [bib38]. However, current studies on data imputation only consider the impact of known data information on model prediction outcomes, neglecting the influence of data label information on imputation prediction research. Furthermore, in industrial applications, data label information also suffers from missing issues. For example, the PhysioBank archives store over 40GB of cardiac medical data, which includes some missing data and a limited amount of data label information [bib23]. Additionally, for legal reasons, hospitals retain a large volume of unlabeled cardiac data. This necessitates that data imputation algorithms not only fill in missing values but also consider the utilization of incomplete label information, thereby better serving down-stream data analysis tasks. To address these issues, this paper proposes a novel missing data imputation model based on semi-supervised learning, SEGAN, which utilizes both known data values and a subset of data labels to complete missing data. The SEGAN model primarily consists of three modules: a generator, a discriminator, and a classifier [bib24]. Specifically, the SEGAN model constructs a classifier based on semi-supervised learning, which drives the generator to complete missing values according to known data and data labels, and outputs the completed data. The discriminator aims to distinguish as much as possible between genuine data and the data completed by the generator. In addition, the SEGAN model introduces a hint matrix to the discriminator, encoding the missing status of part of the data. The main contributions of the paper can be summarized as follows:
@@
(1) We propose a missing data imputation model based on semi-supervised learning, SEGAN, which adopts a semi-supervised learning strategy to correct the completion of incomplete data with partial labels. To the best of our knowledge, this model is the first to utilize data label information for data imputation studies.
@@
(2) Within the SEGAN model, the semi-supervised learning classifier is capable of inferring labels for unlabeled data, thereby requiring the generator to estimate or fill in missing data using existing data and data labels. Meanwhile, the model enhances the discriminator’s accuracy in comparing known data with completed data by using the hint matrix.
@@
(3) We provide theoretical proof that the SEGAN model, based on the missing hint matrix, which can effectively understand information distribution of the incomplete data.
(4) We conducted extensive experiments in this research to demonstrate that the SEGAN model significantly outperforms other existing data imputation methods.
@@
The completion of incomplete structured data essentially requires predicting missing data based on known information, enhancing the integrity of structured data to prevent missing data issues from affecting subsequent decision-making analyses negatively. The simplest method to address missing data is to directly remove samples with missing data from the dataset. However, this approach is only suitable when the volume of incomplete data is small, and the removal of data can be ensured not to affect the analysis results  [bib28]. Such method can easily reduce the information transmitted by the data and severely impact the outcomes of subsequent analysis. Therefore, to mitigate the problems caused by missing data, experts and scholars have focused on research into data imputation methods  [bib29] [bib30] [bib31] [bib32]  [bib33]  [bib34]  [bib35]  [bib36]. Their goal is to predict missing data based on known information, enhance data integrity, and ensuring reliable and robust analytical results. Incomplete structured data completion methods can be categorized into those based on statistical techniques, machine learning-based methods, and deep learning-based methods, depending on the type of prediction model used. More specifically, statistical techniques for data imputation can be categorized based on different statistical mechanisms into methods that rely on statistical information and those based on similarity [bib27]. Methods that use statistical information, such as mean, median, or mode imputation algorithms [bib2], estimate missing values by utilizing statistical data descriptors. Additionally, similarity-based imputation algorithms predict missing data for a target sample using the mean of known values from multiple similar samples. This category includes algorithms such as the K-Nearest Neighbor Imputation (KNNI) [bib3] and Cold Deck Imputation (CDI) [bib4]. Machine learning based imputation methods involve training one or more predictive models to estimate and fill in missing values. Depending on the machine learning model utilized, these methods can be classified into three types: decision tree-based, linear regression-based, and data compression-based imputation algorithms [bib25]. Decision tree-based imputation algorithms construct a predictive decision tree model for each feature with missing data. Examples include the XGBoost Imputation (XGBI) [bib5] algorithm, which leverages the XGBoost framework, and the MissForest Imputation (Miss-FI) [bib6] algorithm, which is based on random forests. Linear regression-based imputation algorithms build multiple linear regression models for each feature with incomplete data, for instance the Multivariate Imputation by Chained Equations (MICE)  [bib7] algorithm, which employs a series of chained equations, and the Imputation via Individual Model (IIM) [bib8], which constructs individual regression models for each feature. Data compression-based imputation algorithms create a data compression predictive model for the entire set of missing data, predicting missing values using singular value decomposition and data reconstruction strategies. This includes the Soft-impute (SI) [bib9] algorithm, which applies a soft-thresholding approach, the Matrix Factorization Imputation (MFI)  [bib10], which factors the data matrix, and the Principal Component Analysis Imputation (PCAI) [bib11], which uses PCA for data imputation. Deep learning-based imputation methods vary according to the specific deep learning models employed and primarily include algorithms based on multi-layer perceptrons, autoencoders, and generative adversarial networks [bib24]. Multi-layer perceptron-based imputation algorithms construct a predictive model using a multi-layer perceptron (MLP) for each incomplete feature. This includes the Multi-Layer Perceptron Imputation (MLPI) [bib12] and the Round-Robin Sinkhorn Imputation (RRSI) [bib13] algorithms. Autoencoder-based imputation algorithms compress incomplete input data into a shallow vector, which is then reconstructed by the decoder into a matrix that closely resembles the input data matrix. Autoencoder-based imputation methods include the Multiple Imputation Denoising Auto-Encoder (MIDAE) [bib14], the Variational Auto-Encoder Imputation (VAEI) [bib15], the Heterogeneous-Incomplete VAE (HIVAE) [bib16], and the Missing Data Importance-Weighted Auto-Encoder Model (MIWAE) [bib17]. Generative adversarial network-based imputation methods design an imputation module to generate missing data that closely approximates the true known distribution and construct a discriminator module to differentiate between imputed and genuine data as accurately as possible [bib26]. This category includes the Generative Adversarial Imputation Network (GAIN) [bib18], the Graph Imputation Neural Network (GINN)  [bib19], the Efficient and Effective Data Imputation System with Influence Functions (EDIT) [bib20], and the Scalable Imputation System (SCIS) [bib21], which estimates the number of samples for generative adversarial imputation. In summary, existing imputation methods for incomplete structured data focus solely on data completion and its processing, neglecting the impact of the imputed results on downstream decision-making outcomes, which leads to lack of effective support of subsequent decision analysis [bib22].
@@ Methods
Fig 1: SEGAN model The architecture of the missing data imputation model SEGAN.
@@
Furthermore, the SEGAN model integrates multi-layer perceptrons (MLPs) with its three primary modules. An MLP is a classical neural network model comprising multiple layers of neurons. Given a multivariate data sample $\mathbf{X}$, the MLP unit outputs a reconstructed multivariate data matrix $\overline{X}$. The update steps for the MLP unit are as follows:
@@
This section will discuss the three modules of the SEGAN model and provide a theoretical prove of the SEGAN model based on the hint matrix and the classifier.
@@
The SEGAN model adopts the traditional concept of generative adversarial models [bib24], motivating the generator to learn the genuine data distribution through the optimization process of the generator and discriminator. Compared to a standard discriminator, the discriminator in the SEGAN model produces a probability matrix, where each probability value indicates the likelihood that each value in the multivariate data $\hat{\mathbf{X}}$, completed after imputation, is genuine. The goal of the discriminator is to differentiate as much as possible between the genuine data and the generated data in $\hat{\mathbf{X}}$. Inspired by the GAIN algorithm [bib6], the SEGAN model introduces a concept known as the ”missingness hint matrix” $\mathbf{R}$, which stores information about the missing status of certain data parts and can be defined as follows:
@@
$\mathbf{m}_{i}=\mathbf{W}_{x}\mathbf{x}_{i}+\mathbf{b}_{x}\quad$ (6)
@@
Here, $\hat{\mathbf{X}}$ represents the multivariate time series data samples filled in by the generator $\mathbf{G}$, $D\left(\hat{\mathbf{X}},\mathbf{R}\right)$ is the discrimination result from discriminator $\mathbf{D}$, and $C\left(\hat{\mathbf{X}}\right)$ is the prediction result from classifier $\mathbf{C}$. Other notations such as $\mathbf{M}$ denote the mask matrix for the incomplete data $\mathbf{X}$, $L_{CE}$ is the cross-entropy loss function, $\mathbf{y}$ is the true label of $\hat{\mathbf{X}}$, and $\beta$ represents a hyperparameter.
@@
According to minimax game theory, the SEGAN model, which incorporates a missingness hint matrix and a classifier, will have its generated data distribution and predicted label distribution converge to the true data distribution and true label distribution upon reaching Nash equilibrium. Firstly, Theorem 1 demonstrates that the missingness hint matrix can aid in determining the equilibrium points of the USEGAN model $V(G,D)$. Subsequently, Theorem 2 proves that when converging to Nash equilibrium, the data distribution generated by $G$ and the label distribution predicted by $C$ will each coincide with the true data distribution and true label distribution, respectively. It is noteworthy that the theorems below are predicated on the assumption of relative independence between $\mathbf{M}$ and $\mathbf{X}$, meaning that the multivariate data is Missing Completely At Random (MCAR) [12].
@@
For a given missingness hint matrix $\mathbf{R}$, the equilibrium state of the USEGAN model $V(G,D)$ can be uniquely determined. At this equilibrium, the data distribution $p$ generated by the USEGAN model and the true data distribution $\hat{p}$ are the same under the given conditions, that is, $p(\mathbf{X}|\mathbf{M})=p(\mathbf{X}|\mathbf{1})$. The proof of Theorem 1 consists of three parts and extends the research of the hint mechanism [6] in the imputation of multivariate data: First, it establishes the existence of an optimal discriminator $\mathbf{D}^{*}$ for the USEGAN model $V(G,D)$ when the generator $G$ is fixed; second, it demonstrates that a discriminator $\mathbf{D}^{*}$ without the missingness hint matrix $\mathbf{R}$ would lead to $G$ generating multiple data distributions; and finally, it proves that the missingness hint matrix $\mathbf{R}$ ensures that $G$ generates a unique true data distribution.
@@
It is important to emphasize that when $\mathbf{R}$ and $M$ are independent, Equation (17) can be rewritten as $\hat{p}(\mathbf{X}|m_{i}^{j}=e)=\hat{p}(\mathbf{X})$. Here, $\hat{p}(\mathbf{X}|m_{i}^{j}=e)=\hat{p}(\mathbf{X})$, and $e\in\{0,1\}$. This relationship holds true only under the following condition:
@@
Given that $\hat{p}(\mathbf{X}|\mathbf{1})$ represents the true and unique data distribution, it follows that any distribution that satisfies Equation (17) must be unique. In other words, the equilibrium position of the USEGAN model $V(G,D)$, which depends on $\mathbf{R}$, can be uniquely determined. At this point, the data distribution $\hat{p}(\mathbf{X}|\mathbf{M})$ generated by the generator $G$ will be consistent with the actual data distribution.
@@
The proof is as follows: In the SEGAN model $U(G,D,C)=V(G,D)+\beta L(G,C)$, the term $L(G,C)$ can be defined as:
@@
Here, the second equation arises from the definition of the cross-entropy loss function, that is, $L_{CE}(q,q^{\prime})=-\sum_{x}q(x)\log q^{\prime}(x)$. This means that for the SEGAN model, minimizing $L(G,C)$ is equivalent to minimizing the Kullback-Leibler divergence $D_{KL}(p_{l}(\cdot|\hat{X})\|p_{c}(\cdot|\hat{X}))$. The latter is non-negative and can only be zero when $p_{l}(\cdot|\hat{X})=p_{c}(\cdot|\hat{X})$. Furthermore, according to Theorem 1, it can be seen that when the model reaches equilibrium, $\hat{p}(\mathbf{X}|\mathbf{M})=\hat{p}(\mathbf{X}|\mathbf{1})$. Therefore, it is only when the SEGAN model reaches Nash equilibrium that $\hat{p}(\mathbf{X}|\mathbf{M})=\hat{p}(\mathbf{X}|\mathbf{1})$ and $p_{l}(\cdot|\hat{X})=p_{c}(\cdot|\hat{X})$, meaning that the data distribution created by generator $G$ and the label distribution predicted by classifier $C$ will align with the actual data distribution and label distribution.
@@
This This section evaluates and compares the imputation performance of the proposed SEGAN algorithm with 15 existing completion models. All algorithms are implemented in the Python environment, and all experiments are conducted on an Intel Core server with a 2.80GHz processor, equipped with a TITAN Xp 12GiB GPU and 192GB RAM.
@@
To thoroughly verify the effectiveness of the SEGAN algorithm under real and complex missing data scenarios, three real-world incomplete datasets were used in the experiments of this chapter. (1) The COVID-19 Trial Tracker (Tria) dataset  [bib41] presents clinical trial data on COVID-19 research and tracks the availability of study outcomes. The Tria dataset records 9 features of 6,433 trials and contains a 9.63% missing data rate. (2) The Emergency Declarations Timeline (Emer) dataset  [bib40] chronicles the emergency declarations and mitigation policies of various U.S. states starting from January 20, 2020, comprising 8,364 samples with 22 features and a 62.69% missing data rate. (3) The Government Response Timeline (Resp) dataset  [bib39] compiles government responses starting from January 1, 2020, including 200,737 samples with 19 features, among which 5.66% of the data is missing.
Metrices
@@
In the experiments, Root Mean Squared Error (RMSE) is employed to measure the imputation accuracy of the completion models. A lower RMSE value indicates better model performance. Additionally, 20% of the known data is randomly deleted during the experiments, and the RMSE is calculated based on these data. Each evaluation metric value refers to the average of the results from five experiments under different random data splits. Moreover, in the experiments, min-max normalization is applied to numerical features on each dataset to prevent certain dimensions from dominating the model training, and one-hot encoding is used to represent categorical features.
Baseline
@@
The experiments include 15 existing state-of-the-art imputation methods for comparison, which comprise 5 machine learning-based imputation methods: MissForest (MissF), XGBoost Imputer (XGBI), Bayesian PCA (Baran), Multiple Imputation by Chained Equations (MICE), and Iterative Collaborative Filtering for Recommender Systems (ICLR); 2 imputation methods based on Multilayer Perceptrons: DataWig and Robust Recursive Imputation (RRSI); 5 imputation methods based on autoencoders: Multiple Imputation using Denoising Autoencoders (MIDAE), Variational Autoencoder Imputer (VAEI), Multiple Imputation with Wasserstein Autoencoders (MIWAE), Efficient Data Imputation with Decision trees (EDDI), and Heteroscedastic Imputation Variational Autoencoder (HIVAE); 2 imputation models based on Generative Adversarial Networks: Graph Imputation Neural Network (GINN) and Generative Adversarial Imputation Nets (GAIN); and the imputation algorithm optimization framework EDIT.
Hyper-parameters

### v3

# SEGAN: A Semi-Supervised Learning Method for Missing Data Imputation

@@ Abstract
In many practical real-world applications, data missing is a very common phenomenon, making the development of data-driven artificial intelligence theory and technology increasingly difficult. Data imputation is an important method for missing data preprocessing. Most existing missing data imputation models directly use the known information in the missing data set but ignore the impact of the data label information contained in the data set on the missing data imputation model. In this paper, we propose a missing data imputation model based on semi-supervised learning, named SEGAN, which mainly includes three important modules: generator, discriminator, and classifier. In the SEGAN model, the classifier enables the generator to make more full use of known data and its label information when predicting missing data values. In addition, the SEGAN model introduces a missing hint matrix to allow the discriminator to more effectively distinguish between known data and data filled by the generator. This paper theoretically proves that the SEGAN model that introduces a classifier and a missing hint matrix can learn the real known data distribution characteristics when reaching Nash equilibrium. Extensive experiments on three public real-world datasets demonstrate that, SEGAN yields a more than 10% accuracy gain, compared with the state-of-the-art approaches.
@@ Introduction
In the information age of today, the issue of data quality has become increasingly pivotal, with the problem of missing data being particularly widespread. This issue is not confined to any single sector but is prevalent across various industries such as healthcare, meteorology, and advertising, and it has a significant impact on data analysis and decision-making. Missing data can lead to incompleteness of information and may also cause biases in data analysis, thereby affecting the accuracy and reliability of decision-making. Faced with this challenge, the academic and industrial communities have devoted efforts to researching data imputation techniques, aiming to find more effective methods to address this issue [bib1].
Historically, researchers have developed a variety of data imputation techniques to address the issue of missing data [bib44] [bib45] [bib46] [bib47] [bib48] [bib49] [bib50]. Early methods often relied on statistics, for instance filling in missing values by inserting mean, median, or mode [bib2] [bib3]. While these methods are straightforward in computation, they overlook the potential correlations and dynamic changes among data, often failing to achieve the desired imputation result. With the development of machine learning and deep learning methods, parametric imputation algorithms based on these methods have emerged [bib1]. These algorithms predict missing values by learning the complex relationships between data, demonstrating more accurate and efficient imputation capabilities.
Moreover, the information contained in data labels is crucial for research on data imputation and data analysis tasks [bib37]  [bib38]. However, current studies on data imputation only consider the impact of known data information on model prediction outcomes, neglecting the influence of data label information on imputation prediction research. Furthermore, in industrial applications, data label information also suffers from missing issues. For example, the PhysioBank archives store over 40GB of cardiac medical data, which includes some missing data and a limited amount of data label information [bib23]. Additionally, for legal reasons, hospitals retain a large volume of unlabeled cardiac data. This necessitates that data imputation algorithms not only fill in missing values but also consider the utilization of incomplete label information, thereby better serving downstream data analysis tasks.
To address these issues, this paper proposes a novel missing data imputation model based on semi-supervised learning, SEGAN, which utilizes both known data values and a subset of data labels to complete missing data. The SEGAN model primarily consists of three modules: a generator, a discriminator, and a classifier [bib24]. Specifically, the SEGAN model constructs a classifier based on semi-supervised learning, which drives the generator to complete missing values according to known data and data labels, and outputs the completed data. The discriminator aims to distinguish as much as possible between genuine data and the data completed by the generator. In addition, the SEGAN model introduces a hint matrix to the discriminator, encoding the missing status of part of the data. The main contributions are summarized as follows.
@@
  • We propose a missing data imputation model based on semi-supervised learning, SEGAN, which adopts a semi-supervised learning strategy to correct the imputation of incomplete data with partial labels. To the best of our knowledge, this model is the first to utilize data label information for data imputation studies.
@@
  • Within the SEGAN model, the semi-supervised learning classifier is capable of inferring labels for unlabeled data, thereby requiring the generator to estimate or fill in missing data using existing data and data labels. Meanwhile, the model enhances the discriminator’s accuracy in comparing known data with completed data by using the hint matrix.
@@
  • We provide theoretical proof that the SEGAN model, based on the missing hint matrix, can effectively understand the information distribution of incomplete data.
  • Extensive experiments using several real-life datasets demonstrate the remarkable performance enhancement of SEGAN, compared with the state-of-the-art approaches.
The code is shared openly on Github (https://github.com/niunaicoke/SEGAN-A-Semi-Supervised-Learning-Method-for-Missing-Data-Imputation).
@@
The imputation of incomplete structured data essentially requires predicting missing data based on known information, enhancing the integrity of structured data to prevent missing data issues from affecting subsequent decision-making analyses negatively. The simplest method to address missing data is to directly remove samples with missing data from the dataset. However, this approach is only suitable when the volume of incomplete data is small, and the removal of data can be ensured not to affect the analysis results  [bib28]. Such a method can easily reduce the information transmitted by the data and severely impact the outcomes of subsequent analysis. Therefore, to mitigate the problems caused by missing data, experts and scholars have focused on research into data imputation methods  [bib29] [bib30] [bib31] [bib32]  [bib33]  [bib34]  [bib35]  [bib36]. Their goal is to predict missing data based on known information, enhance data integrity, and ensure reliable and robust analytical results. Incomplete structured data imputation methods can be categorized into those based on statistical techniques, machine learning-based methods, and deep learning-based methods, depending on the type of prediction model used.
More specifically, statistical techniques for data imputation can be categorized based on different statistical mechanisms into methods that rely on statistical information and those based on similarity [bib27]  [bib43] [bib42]. Methods that use statistical information, such as mean, median, or mode imputation algorithms [bib2], estimate missing values by utilizing statistical data descriptors. Additionally, similarity-based imputation algorithms predict missing data for a target sample using the mean of known values from multiple similar samples. This category includes algorithms such as the K-Nearest Neighbor Imputation (KNNI) [bib3] and Cold Deck Imputation (CDI) [bib4].
Machine learning based imputation methods involve training one or more predictive models to estimate and fill in missing values. Depending on the machine learning model utilized, these methods can be classified into three types: decision tree-based, linear regression-based, and data compression-based imputation algorithms [bib25]. Decision tree-based imputation algorithms construct a predictive decision tree model for each feature with missing data. Examples include the XGBoost Imputation (XGBI) [bib5] algorithm, which leverages the XGBoost framework, and the MissForest Imputation (Miss-FI) [bib6] algorithm, which is based on random forests. Linear regression-based imputation algorithms build multiple linear regression models for each feature with incomplete data, for instance, the Multivariate Imputation by Chained Equations (MICE)  [bib7] algorithm, which employs a series of chained equations, and the Imputation via Individual Model (IIM) [bib8], which constructs individual regression models for each feature. Data compression-based imputation algorithms create a data compression predictive model for the entire set of missing data, predicting missing values using singular value decomposition and data reconstruction strategies. This includes the Soft-impute (SI) [bib9] algorithm, which applies a soft-thresholding approach, the Matrix Factorization Imputation (MFI)  [bib10], which factors the data matrix, and the Principal Component Analysis Imputation (PCAI) [bib11], which uses PCA for data imputation.
Deep learning-based imputation methods vary according to the specific deep learning models employed and primarily include algorithms based on multi-layer perceptrons, autoencoders, and generative adversarial networks [bib24]. Multi-layer perceptron-based imputation algorithms construct a predictive model using a multi-layer perceptron (MLP) for each incomplete feature. This includes the Multi-Layer Perceptron Imputation (MLPI) [bib12] and the Round-Robin Sinkhorn Imputation (RRSI) [bib13] algorithms. Autoencoder-based imputation algorithms compress incomplete input data into a shallow vector, which is then reconstructed by the decoder into a matrix that closely resembles the input data matrix. Autoencoder-based imputation methods include the Multiple Imputation Denoising Auto-Encoder (MIDAE) [bib14], the Variational Auto-Encoder Imputation (VAEI) [bib15], the Heterogeneous-Incomplete VAE (HIVAE) [bib16], and the Missing Data Importance-Weighted Auto-Encoder Model (MIWAE) [bib17]. Generative adversarial network-based imputation methods design an imputation module to generate missing data that closely approximates the true known distribution and construct a discriminator module to differentiate between imputed and genuine data as accurately as possible [bib26]. This category includes the Generative Adversarial Imputation Network (GAIN) [bib18], the Graph Imputation Neural Network (GINN)  [bib19], the Efficient and Effective Data Imputation System with Influence Functions (EDIT) [bib20], and the Scalable Imputation System (SCIS) [bib21], which estimates the number of samples for generative adversarial imputation.
In summary, existing imputation methods for incomplete structured data focus solely on data imputation and its processing, neglecting the impact of the imputed results on downstream decision-making outcomes, which leads to a lack of effective support of subsequent decision analysis [bib22].
@@ Methods
Fig 1: The architecture of SEGAN.
@@
Furthermore, the SEGAN model integrates multi-layer perceptrons (MLPs) with its three primary modules. An MLP is a classical neural network model comprising multiple layers of neurons. Given a multivariate data sample $\mathbf{X}$, the MLP unit outputs a reconstructed data matrix $\overline{X}$. The update steps for the MLP unit are as follows:
@@
This section will discuss the three modules of the SEGAN model and provide theoretical proof of the SEGAN model based on the hint matrix and the classifier.
@@
The SEGAN model adopts the traditional concept of generative adversarial models [bib24], motivating the generator to learn the genuine data distribution through the optimization process of the generator and discriminator. Compared to a standard discriminator, the discriminator in the SEGAN model produces a probability matrix, where each probability value indicates the likelihood that each value in the multivariate data $\hat{\mathbf{X}}$, completed after imputation, is genuine. The goal of the discriminator is to differentiate as much as possible between the genuine data and the generated data in $\hat{\mathbf{X}}$. Inspired by the GAIN algorithm [bib6], the SEGAN model introduces a concept known as the “missingness hint matrix” $\mathbf{R}$, which stores information about the missing status of certain data parts and can be defined as follows:
@@
$\mathbf{m}_{i}=\mathbf{W}_{x}\mathbf{x}_{i}+\mathbf{b}_{x}$ (6)
@@
Here, $\hat{\mathbf{X}}$ represents the multivariate time series data samples filled in by the generator $\mathbf{G}$, $D(\hat{\mathbf{X}},\mathbf{R})$ is the discrimination result from discriminator $\mathbf{D}$, and $C(\hat{\mathbf{X}})$ is the prediction result from classifier $\mathbf{C}$. Other notations such as $\mathbf{M}$ denote the mask matrix for the incomplete data $\mathbf{X}$, $L_{CE}$ is the cross-entropy loss function, $\mathbf{y}$ is the true label of $\hat{\mathbf{X}}$, and $\beta$ represents a hyperparameter.
@@
According to minimax game theory, the SEGAN model, which incorporates a missingness hint matrix and a classifier, will have its generated data distribution and predicted label distribution converges to the true data distribution and true label distribution upon reaching Nash equilibrium. Firstly, Theorem 1 demonstrates that the missingness hint matrix can aid in determining the equilibrium points of the USEGAN model $V(G,D)$. Subsequently, Theorem 2 proves that when converging to Nash equilibrium, the data distribution generated by $G$ and the label distribution predicted by $C$ will each coincide with the true data distribution and true label distribution, respectively. It is noteworthy that the theorems below are predicated on the assumption of relative independence between $\mathbf{M}$ and $\mathbf{X}$, meaning that the multivariate data is Missing Completely At Random (MCAR) [bib25].
@@
For a given missingness hint matrix $\mathbf{R}$, the equilibrium state of the USEGAN model $V(G,D)$ can be uniquely determined. At this equilibrium, the data distribution $p$ generated by the USEGAN model and the true data distribution $\hat{p}$ are the same under the given conditions, that is, $p(\mathbf{X}|\mathbf{M})=p(\mathbf{X}|\mathbf{1})$.
Proof. The proof of Theorem 1 consists of three parts and extends the research of the hint mechanism [6] in the imputation of multivariate data: First, it establishes the existence of an optimal discriminator $\mathbf{D}^{*}$ for the USEGAN model $V(G,D)$ when the generator $G$ is fixed; second, it demonstrates that a discriminator $\mathbf{D}^{*}$ without the missingness hint matrix $\mathbf{R}$ would lead to $G$ generating multiple data distributions; and finally, it proves that the missingness hint matrix $\mathbf{R}$ ensures that $G$ generates a unique true data distribution.
@@
It is important to emphasize that when $\mathbf{R}$ and $M$ are independent, Equation (17) can be rewritten as $\hat{p}(\mathbf{X}|m_{i}^{j}=e)=\hat{p}(\mathbf{X})$. Here, $\hat{p}(\mathbf{X}|m_{i}^{j}=e)=\hat{p}(\mathbf{X})$, and $e\in\{0,1\}$. This relationship holds true only under the following conditions:
@@
Given that $\hat{p}(\mathbf{X}|\mathbf{1})$ represents the true and unique data distribution, it follows that any distribution that satisfies Equation (17) must be unique. In other words, the equilibrium position of the USEGAN model $V(G,D)$, which depends on $\mathbf{R}$, can be uniquely determined. At this point, the data distribution $\hat{p}(\mathbf{X}|\mathbf{M})$ generated by the generator $G$ will be consistent with the actual data distribution. $\square$
@@
Proof. In the SEGAN model $U(G,D,C)=V(G,D)+\beta L(G,C)$, the term $L(G,C)$ can be defined as:
@@
Here, the second equation arises from the definition of the cross-entropy loss function, that is, $L_{CE}(q,q^{\prime})=-\sum_{x}q(x)\log q^{\prime}(x)$. This means that for the SEGAN model, minimizing $L(G,C)$ is equivalent to minimizing the Kullback-Leibler divergence $D_{KL}(p_{l}(\cdot|\hat{X})\|p_{c}(\cdot|\hat{X}))$. The latter is non-negative and can only be zero when $p_{l}(\cdot|\hat{X})=p_{c}(\cdot|\hat{X})$. Furthermore, according to Theorem 1, it can be seen that when the model reaches equilibrium, $\hat{p}(\mathbf{X}|\mathbf{M})=\hat{p}(\mathbf{X}|\mathbf{1})$. Therefore, it is only when the SEGAN model reaches Nash equilibrium that $\hat{p}(\mathbf{X}|\mathbf{M})=\hat{p}(\mathbf{X}|\mathbf{1})$ and $p_{l}(\cdot|\hat{X})=p_{c}(\cdot|\hat{X})$, meaning that the data distribution created by generator $G$ and the label distribution predicted by classifier $C$ will align with the actual data distribution and label distribution. $\square$
@@
This section evaluates and compares the imputation performance of the proposed SEGAN algorithm with 12 existing imputation models. All algorithms are implemented in the Python environment, and all experiments are conducted on an Intel Core server with a 2.80GHz processor, equipped with a TITAN Xp 12GiB GPU and 192GB RAM.
@@
Dataset. To thoroughly verify the effectiveness of the SEGAN algorithm under real and complex missing data scenarios, three real-world incomplete datasets were used in the experiments of this chapter. (1) The COVID-19 Trial Tracker (Tria) dataset  [bib41] presents clinical trial data on COVID-19 research and tracks the availability of study outcomes. The Tria dataset records 9 features of 6,433 trials and contains a 9.63% missing data rate. (2) The Emergency Declarations Timeline (Emer) dataset  [bib40] chronicles the emergency declarations and mitigation policies of various U.S. states starting from January 20, 2020, comprising 8,364 samples with 22 features and a 62.69% missing data rate. (3) The Government Response Timeline (Resp) dataset  [bib39] compiles government responses starting from January 1, 2020, including 200,737 samples with 19 features, among which 5.66% of the data is missing.
@@
Metrices. In the experiments, Root Mean Squared Error (RMSE) is employed to measure the imputation accuracy of the imputation models. A lower RMSE value indicates better model performance. Additionally, 20% of the known data is randomly deleted during the experiments, and the RMSE is calculated based on these data. Each evaluation metric value refers to the average of the results from five experiments under different random data splits. Moreover, in the experiments, min-max normalization is applied to numerical features on each dataset to prevent certain dimensions from dominating the model training, and one-hot encoding is used to represent categorical features.
@@
Baseline. The experiments include 12 existing state-of-the-art imputation methods for comparison, which comprise 5 machine learning-based imputation methods: MissForest (MissF), XGBoost Imputer (XGBI), Bayesian PCA (Baran), Multiple Imputation by Chained Equations (MICE), and Iterative Collaborative Filtering for Recommender Systems (ICLR); 2 imputation methods based on Multilayer Perceptrons: DataWig and Robust Recursive Imputation (RRSI); 5 imputation methods based on autoencoders: Multiple Imputation using Denoising Autoencoders (MIDAE), Variational Autoencoder Imputer (VAEI), Multiple Imputation with Wasserstein Autoencoders (MIWAE), Efficient Data Imputation with Decision trees (EDDI), and Heteroscedastic Imputation Variational Autoencoder (HIVAE); 2 imputation models based on Generative Adversarial Networks: Graph Imputation Neural Network (GINN) and Generative Adversarial Imputation Nets (GAIN); and the imputation algorithm optimization framework EDIT.